# Asymptotic regularity of a generalised stochastic Halpern scheme

### v2

@@ 2. Preliminaries and basic lemmas
[…]
@@
We note that whenever $\Phi$ is a rate of convergence for $X_{n}\to 0$ almost surely, then for every $\varepsilon>0$, $\Phi(\varepsilon,\cdot)$ is a rate of convergence for $\mathbb{P}\left(\sup_{n\geq N}\,(X_{n}\geq\varepsilon)\right)\to 0$ in the nonstochastic sense.
@@
for all $n\in\mathbb{N}$ where $\sum_{n=0}^{\infty}a_{n}=\infty$, lim $\limsup b_{n}\leq 0$ and $\sum_{n=0}^{\infty}c_{n}<\infty$. Then $s_{n}\to 0$.
@@
for all $n\in\mathbb{N}$ and so the rate for $\mathbb{E}[X_{n}]\to 0$ follows by Lemma 2.2. For the rate for the almost sure convergence, observe that $\chi(\lambda\varepsilon/4)\leq\varphi_{K,\theta,\chi}(\lambda\varepsilon/2)$ as $\theta(k,b)\geq k$. Hence, one can proceed as in the proof of Lemma 2.3 to show that
@@
for any $N$ where, using $\chi(\lambda\varepsilon/4)\leq\varphi_{K,\theta,\chi}(\lambda\varepsilon/2)$, we then can conclude $\mathbb{P}\left(\exists n\geq N(U_{n}\geq\varepsilon)\right)<\lambda$ for $N=\varphi_{K,\theta,\chi}(\lambda\varepsilon/2)$. ∎
@@ 3.1. Basic results and rates of asymptotic regularity
We begin with some fundamental recursive inequalities for the iterations generated by the iteration (1):
@@
Let $(x_{n}),(y_{n})$ be the sequences generated by (1). Then the following recurrence relations hold pointwise everywhere for all $n\in\mathbb{N}$:
@@
We now move to our first quantitative result which presents a rate of asymptotic regularity for the sequence $(x_{n})$, both in expectation and in probability. For that we introduce a first central assumption on the boundedness of the iteration (1) in expectation, as commonly made in the literature (see e.g. hypothesis $(\mathrm{H}_{1})$ in [5] of which this assumption here is a natural extension to the generalised iteration (1)):
@@
Let $(x_{n}),(y_{n})$ be the sequences generated by (1). Suppose that $\sum_{n=0}^{\infty}\alpha_{n}=\infty$ with rate of divergence $\theta$, that
@@ 3.2. Asymptotic regularity relative to the mappings
We now move on to establishing rates of asymptotic regularity for the iterations relative to the mappings. For that, we will actually see a crucial dichotomy, where results based on the use of just one of the mappings $U$ or $T$ are comparatively straightforward, whereas for the general case where neither $U$ nor $T$ trivialize, we rely on a geometric argument for establishing a rate of asymptotic regularity relative to $U$ for the sequence $(y_{n})$, which requires both a uniform convexity of the space $X$ along with a uniform integrability assumption on $(\left\lVert Uy_{n}-y_{n}\right\rVert)$. As such, before we move on to these results, we first give rates of asymptotic regularity relative to the mappings for the remaining cases, dependent on the relevant rates for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$:
@@
Proof.
By taking the expectation, the rates immediately follow. ∎
@@
In the special case of $U:=\mathrm{Id}$ and $\delta_{n}:=0$, the iteration (1) collapses to the stochastic Halpern iteration (1). We then have trivial rates for $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ and $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely and so, in that case, we get under the assumptions of Theorems 3.2 and 3.4 (and also under suitable monotonicity assumptions of the rates involved) that
@@
with $\varphi_{K,\theta,\chi}$ defined as in Theorem 3.2 and $\rho_{1},\rho_{3}$ as in Theorem 3.4. Note that this generalises known rates in this case [5, Theorem 3.3], which apply only to specific choices of the parameters. In a similar way, we get a new rate for $\left\lVert Tx_{n}-x_{n}\right\rVert\to 0$ almost surely, though we do not spell it out here.
In the special case of $T:=\mathrm{Id}$ and $\xi_{n}:=0$, the iteration (1) collapses to a stochastic variant of the Krasnoselskii-Mann iteration with Tikhonov regularization terms (1) (and even a slight extension by allowing general anchors $u$). In that case, we do not need to rely on the geometric arguments discussed in the next part of this section and can instead directly derive rates of convergence for $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ and $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely, essentially following the approach of [40] (see also [7]).
@@
Let $(x_{n}),(y_{n})$ be the sequences generated by (1) where $T=\mathrm{Id}$ and $\xi_{n}=0$ for all $n\in\mathbb{N}$. Then the following recurrence relation holds everywhere on $\Omega$ for all $n\in\mathbb{N}$:
@@
Proof.
Given $n\in\mathbb{N}$, we have
pointwise everywhere. ∎
@@
Let $(x_{n}),(y_{n})$ be the sequences generated by (1) where $T=\mathrm{Id}$ and $\xi_{n}=0$ for all $n\in\mathbb{N}$. Also, let $\Lambda>0$ be such that $\Lambda\leq\beta_{n}\leq 1-\Lambda$ for all $n\in\mathbb{N}$. If $\mathbb{E}[\left\lVert y_{n}-y_{n+1}\right\rVert]\to 0$ with rate $\varphi$, $\alpha_{n}\to 0$ with rate $\rho$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\to 0$ with rate $\chi$, then $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ with rate
@@ 3.2.2. The general case
[…]
@@
We now move on to the asymptotic regularity relative to $U$ of the sequence $(y_{n})$. For that, we initially establish $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely using geometric properties of the underlying space and then use quantitative uniform integrability assumption for this random variables to recover the asymptotic regularity relative to $U$ of $(y_{n})$ in expectation. First, we will make the following geometric assumption on the underlying normed space $(X,\left\lVert\cdot\right\rVert)$:
@@
The above modulus also applies to closed balls of any radius centered at any point in the space and for arbitrary convex combinations:
@@
$A_{\lambda,\varepsilon}\subseteq\{\exists n\geq\Gamma(\lambda,\varepsilon)\,(V_{n}\geq\widehat{\varepsilon})\}.$
@@
$\leq\mathbb{P}(\exists n\geq\Gamma(\lambda,\varepsilon)\,(V_{n}\geq\widehat{\varepsilon}))$
@@
Using a slightly different argument first devised in [19, Theorem 3.4] (see also [26, Remark 15] or [30, Remark 3.7] for similar remarks in the context of nonlinear spaces), we can slightly optimize the above rate in the context of moduli of uniform convexity of a special form: Let $\eta(\varepsilon)=\varepsilon\cdot\tilde{\eta}(\varepsilon)$ where $\tilde{\eta}$ is increasing. Then above rate $\Gamma$ holds even with $\widehat{\varepsilon}:=\varepsilon\cdot\Lambda^{2}\cdot\tilde{\eta}(\varepsilon/K^{\prime})$.
@@
Before moving on to the case of fast rates for special choices of scalars and errors, we briefly discuss some natural conditions under which the previous additional assumption of a modulus of uniform integrability for the sequence $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ can be obtained (cf. [15, Chapter 5.4]), and indicate how a modulus arises from these.
A first common assumption to establish uniform integrability is to assume a uniform finite $p$-th moment for $p>1$. Now, in terms of an upper bound $K$ witnessing that assumption, i.e. $\sup_{n\in\mathbb{N}}\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert^{p}]<K$ for $p>1$, we can actually compute a corresponding modulus $\mu$. To see this, note that for any $\varepsilon>0$, setting $X_{n}:=\left\lVert Uy_{n}-y_{n}\right\rVert$, we have
$\mathbb{E}[|X_{n}|1_{A}]\leq\mathbb{E}[|X_{n}|1_{A\cap(|X_{n}|\leq a)}]+\mathbb{E}[|X_{n}|1_{A\cap(|X_{n}|>a)}]\leq a\mathbb{P}(A)+a^{1-p}\mathbb{E}[|X_{n}|^{p}]\leq a\mathbb{P}(A)+\varepsilon/2$
for $a:=(2K/\varepsilon)^{1/(p-1)}$ and $A\in\mathcal{F}$ and thus a modulus of uniform integrability is given by
$\mu(\varepsilon):=\frac{\varepsilon}{2}\left(\frac{\varepsilon}{2K}\right)^{1/(p-1)}.$
Another growth condition commonly imposed is to assume uniform finite expectation of the random variables under a supercoercive function. Then, in terms of an upper bound $K$ witnessing that property, i.e. $\sup_{n\in\mathbb{N}}\mathbb{E}[g(\left\lVert Uy_{n}-y_{n}\right\rVert)]<K$ where $g:[0,\infty)\to[0,\infty)$ is supercoercive, i.e. $g(x)/x\to\infty$ as $x\to\infty$, with a rate of divergence $\kappa:(0,\infty)\to(0,\infty)$, i.e.
$\forall a>0\,\forall x\geq\kappa(a)\left(\frac{g(x)}{x}\geq a\right),$
a corresponding modulus $\mu$ can again be easily computed: Similarly to above we have
$\mathbb{E}[|X_{n}|1_{A}]\leq\kappa(a)\mathbb{P}(A)+\mathbb{E}[g\left(|X_{n}|\right)]/a\leq\kappa(a)\mathbb{P}(A)+\varepsilon/2$
for $a:=2K/\varepsilon$ and thus a modulus of uniform integrability is now given by $\mu(\varepsilon):=\varepsilon/2\kappa\left(2K/\varepsilon\right)$.
@@
Most importantly, we want to show that the required modulus of uniform integrability can be obtained through quantitative integrability assumptions on the errors. Concretely, note that we have $\left\lVert Uy_{n}-y_{n}\right\rVert\leq Y$ pointwise everywhere by Lemma 3.9, so a modulus of integrability for $Y$ immediately yields a modulus of uniform integrability for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$. In particular, if $u$ and $x_{0}$ are chosen to be constant and $K>0$ is such that $\left\lVert x_{0}-p\right\rVert,\left\lVert u-p\right\rVert<K$, and in addition both $\sum_{i=0}^{\infty}\left\lVert\xi_{i}\right\rVert$ and $\sum_{i=0}^{\infty}\left\lVert\delta_{i}\right\rVert$ are integrable with moduli of absolute continuity $\mu_{1}$ and $\mu_{2}$ respectively, by Lemma 3.9 we have
$\mathbb{E}[|X_{n}|1_{A}]\leq 2K\mathbb{P}(A)+2\left(\sum_{i=0}^{\infty}\left(\mathbb{E}[\left\lVert\xi_{i}\right\rVert 1_{A}]+\mathbb{E}[\left\lVert\delta_{i}\right\rVert 1_{A}]\right)\right)$
and thus a modulus of uniform integrability is now given by
$\mu(\varepsilon):=\min\left\{\frac{\varepsilon}{4K},\mu_{1}\left(\frac{\varepsilon}{8}\right),\mu_{2}\left(\frac{\varepsilon}{8}\right)\right\}.$
In this way, we demonstrate that it is possible to shift the integrability condition on $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ to appropriate integrability conditions on the sums of the error terms.

### v3

@@ 2. Preliminaries and basic lemmas
We write $\mathbb{N}^{*}$ for $\mathbb{N}$ without 0. Throughout, if not stated otherwise, we fix an underlying probability space $(\Omega,\mathcal{F},\mathbb{P})$ and all probabilistic notions such as almost sureness refer to that space. Similarly, $X$ will always denote, unless stated otherwise, a normed space with norm $\left\lVert\cdot\right\rVert$. We refer to measurable functions $\Omega\to\mathbb{R}$ as random variables, to measurable functions $\Omega\to X$ as $X$-valued random variables and we refer to sequences of random variables as stochastic processes. In order to ensure that basic properties enjoyed by real-valued random variables are also inherited by $X$-valued random variables, so that in particular our main scheme (1.1) is well-defined, one normally requires some further assumptions on the underlying space (as discussed in detail in [25]). The simplest option is to assume that $X$ is a separable Banach space, though if the reader prefers they can also just assume that $X$ is finite dimensional. Equalities and inequalities involving random variables will always be understood to hold almost surely, even if not explicitly indicated.
@@
We note that whenever $\Phi$ is a rate of convergence for $X_{n}\to 0$ almost surely, then for every $\varepsilon>0$, $\Phi(\varepsilon,\cdot)$ is a rate of convergence for $\mathbb{P}\left(\sup_{n\geq N}\,(X_{n}\geq\varepsilon)\right)\to 0$ as $N\to\infty$ in the nonstochastic sense.
@@
for all $n\in\mathbb{N}$ where $\sum_{n=0}^{\infty}a_{n}=\infty$, $\limsup b_{n}\leq 0$ and $\sum_{n=0}^{\infty}c_{n}<\infty$. Then $s_{n}\to 0$.
@@
for all $n\in\mathbb{N}$ and so the rate for $\mathbb{E}[X_{n}]\to 0$ follows by Lemma $2.$ For the rate for the almost sure convergence, observe that $\chi(\lambda\varepsilon/4)\leq\varphi_{K,\theta,\chi}(\lambda\varepsilon/2)$ as $\theta(k,b)\geq k$. Hence, one can proceed as in the proof of Lemma 2.3 to show that
@@
for any NN where, using $\chi(\lambda\varepsilon/4)\leq\varphi_{K,\theta,\chi}(\lambda\varepsilon/2)$, we then can conclude $\mathbb{P}\left(\exists n\geq N(U_{n}\geq\varepsilon)\right)<\lambda$ for $N=\varphi_{K,\theta,\chi}(\lambda\varepsilon/2)$. ∎
@@ 3.1. Basic results and rates of asymptotic regularity
We begin with some fundamental recursive inequalities for the iterations generated by the iteration (1.1):
@@
Let $(x_{n}),(y_{n})$ be the sequences generated by (1.1). Then the following recurrence relations hold pointwise everywhere for all $n\in\mathbb{N}$:
@@
We now move to our first quantitative result which presents a rate of asymptotic regularity for the sequence $(x_{n})$, both in expectation and in probability. For that we introduce a first central assumption on the boundedness of the iteration (1.1) in expectation, as commonly made in the literature (see e.g. hypothesis $(\mathrm{H}_{1})$ in [4] of which this assumption here is a natural extension to the generalised iteration (1.1)):
@@
Let $(x_{n}),(y_{n})$ be the sequences generated by (1.1). Suppose that $\sum_{n=0}^{\infty}\alpha_{n}=\infty$ with rate of divergence $\theta$, that
@@ 3.2. Asymptotic regularity relative to the mappings
We now move on to establishing rates of asymptotic regularity for the iterations relative to the mappings. For that, we will actually see a crucial dichotomy, where results based on the use of just one of the mappings $U$ or $T$ are comparatively straightforward, whereas for the general case where neither $U$ nor $T$ trivialize, we rely on a geometric argument for establishing a rate of asymptotic regularity relative to $U$ for the sequence $(y_{n})$ which requires a uniform convexity assumption on the space $X$, and in the case of asymptotic regularity in mean further relies on an essential use of the uniform integrability of $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ (which for example immediately derives from the assumption (Hyp), as also discussed in detail later). As such, before we move on to these results, we first give rates of asymptotic regularity relative to the mappings for the remaining cases, dependent on the relevant rates for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$. As all proofs outside this geometric construction are rather routine, they are deferred to the appendix.
@@
In the special case of $U:=\mathrm{Id}$ and $\delta_{n}:=0$, the iteration (1.1) collapses to the stochastic Halpern iteration (1.1). We then have trivial rates for $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ and $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely and so, in that case, we get under the assumptions of Theorems 3.2 and 3.4 (and also under suitable monotonicity assumptions of the rates involved) that
@@
with $\varphi_{K,\theta,\chi}$ defined as in Theorem 3.2 and $\rho_{1},\rho_{3}$ as in Theorem 3.4. Note that this generalises known rates in this case [4, Theorem 3.3], which apply only to specific choices of the parameters. In a similar way, we get a new rate for $\left\lVert Tx_{n}-x_{n}\right\rVert\to 0$ almost surely, though we do not spell it out here.
In the special case of $T:=\mathrm{Id}$ and $\xi_{n}:=0$, the iteration (1.1) collapses to a stochastic variant of the Krasnoselskii-Mann iteration with Tikhonov regularization terms (1.1) (and even a slight extension by allowing general anchors $u$). In that case, we do not need to rely on the geometric arguments discussed in the next part of this section and can instead directly derive rates of convergence for $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ and $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely, essentially following the approach of [45] (see also [6]).
@@
Let $(x_{n}),(y_{n})$ be the sequences generated by (1.1) where $T=\mathrm{Id}$ and $\xi_{n}=0$ for all $n\in\mathbb{N}$. Then the following recurrence relation holds everywhere on $\Omega$ for all $n\in\mathbb{N}$:
@@
Let $(x_{n}),(y_{n})$ be the sequences generated by (1.1) where $T=\mathrm{Id}$ and $\xi_{n}=0$ for all $n\in\mathbb{N}$. Also, let $\Lambda>0$ be such that $\Lambda\leq\beta_{n}\leq 1-\Lambda$ for all $n\in\mathbb{N}$. If $\mathbb{E}[\left\lVert y_{n}-y_{n+1}\right\rVert]\to 0$ with rate $\varphi$, $\alpha_{n}\to 0$ with rate $\rho$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\to 0$ with rate $\chi$, then $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ with rate
@@ 3.2.2. The general case
We now discuss an alternative scenario where, in particular, a random variable $Y$ satisfying (Hyp) can be explicitly constructed if our mappings possess a common fixed point. To be more precise, let us assume that $\mathrm{Fix}T\cap\mathrm{Fix}U\neq\emptyset$ and that $p$ is a common fixed point of $T$ and $U$. Further, instead of making the assumptions (Hyp) or (Hyp), for the rest of this section we now fix
@@
We now move on to the asymptotic regularity relative to $U$ of the sequence $(y_{n})$, which then in conjunction with Theorems 3.4 and 3.5 will allow us to derive all of the other regularity properties relative to mappings. For that, we initially establish $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely using geometric properties of the underlying space. Using that $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ is uniformly integrable, as follows from (Hyp) and hence from Lemma 3.9 in the context of the present assumptions, we can then establish the asymptotic regularity relative to $U$ of $(y_{n})$ in expectation. The quantitative result will then in particular depend on quantitative renderings of both the uniform convexity of the space and the uniform integrability of the sequence. We begin with the geometric assumption on the underlying normed space $(X,\left\lVert\cdot\right\rVert)$:
@@
We will later discuss examples of such moduli of uniform convexity for particular (classes of) spaces (see in particular Lemma 4.10). The above modulus also applies to closed balls of any radius centered at any point in the space and for arbitrary convex combinations:
@@
[…]
@@
$)\displaystyle\leq\mathbb{P}(\exists n\geq\Gamma(\lambda,\varepsilon)\,(V_{n}\geq\widehat{\varepsilon}$))
@@
Using a slightly different argument first devised in [18, Theorem 3.4] (see also [26, Remark 15] or [30, Remark 3.7] for similar remarks in the context of nonlinear spaces), we can slightly optimize the above rate in the context of moduli of uniform convexity of a special form: Let $\eta(\varepsilon)=\varepsilon\cdot\tilde{\eta}(\varepsilon)$ where $\tilde{\eta}$ is increasing. Then the above rate $\Gamma$ holds even with $\widehat{\varepsilon}:=\varepsilon\cdot\Lambda^{2}\cdot\tilde{\eta}(\varepsilon/K^{\prime})$.
@@
By e.g. Lemma 13.1 in [47], such a modulus always exists for integrable $X$. However, it provides a quantitatively different rendering of that property than a simple upper bound on the mean and can in general not be derived from such a bound. We begin with a simple property of such moduli.
@@
A modulus of absolute continuity now motivates the following new notion of a modulus of uniform integrability:
@@
We now obtain the following result on rates of asymptotic regularity in mean, derived from the previous Theorem 3.12 using a modulus of uniform integrability for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$.
@@
The following result collects the qualitative core of the above asymptotic regularity results for (1.1), void of any quantitative considerations, to illustrate the main assumptions more clearly:
Corollary 3.20.
Let $(X,\left\lVert\cdot\right\rVert)$ be uniformly convex and let $(x_{n}),(y_{n})$ be the sequences generated by (1.1) where $\mathrm{Fix}T\cap\mathrm{Fix}U\neq\emptyset$. Assume $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$, $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$ and $\mathbb{E}[\left\lVert x_{0}-p\right\rVert],\mathbb{E}[\left\lVert u-p\right\rVert]<\infty$ together with $\sum_{n=0}^{\infty}\alpha_{n}=\infty$ and $\alpha_{n}\to 0$ as well as
and $\Lambda\leq\beta_{n}\leq 1-\Lambda$ for some $\Lambda>0$. Then $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely and in mean.
To be more precise, Corollary 3.20 follows from Theorems 3.12 and 3.18 in conjunction with Theorem 3.2. As emphasised already, the requirement in Theorem 3.18 that $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ comes equipped with a modulus of uniform integrability disappears in the qualitative result: Under the assumptions of this section, (Hyp) is automatically satisfied, and in particular $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ is dominated by some integrable random variable $Y$ (explicitly definable by Lemma 3.9) and is thus automatically uniformly integrable. The modulus is only required in order to construct our rate of convergence.
Indeed, both Theorem 3.12 and 3.18 provide rather complex constructions for the corresponding rates of asymptotic regularity, featuring an interplay of many different moduli, in particular of the moduli of uniform convexity $\eta$ and uniform integrability $\mu$. While concrete instantiations for the modulus $\eta$ for special (classes of) spaces will be discussed later on (see Lemma 4.10) in the context of standard assumptions on the errors and the parameters which result in particularly fast rates, here we illustrate how a corresponding modulus $\mu$ can be derived in more concrete situations.
We note that under our general assumption (Hyp) whereby $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ is dominated by some $Y$ with finite mean, a modulus of absolute continuity for $Y$ (in the sense of Definition 3.14) is clearly also a modulus of uniform integrability for the sequence $(\left\lVert Uy_{n}-y_{n}\right\rVert)$, and our quantitative uniform integrability requirement is thus reduced to a quantitative continuity property of the bound $Y$. Here the bounds $D,E$ for the summability assumption on the error terms do not suffice to derive such a modulus explicitly, as they do not allow for a construction of corresponding moduli of absolute continuity for the series, but provided these assumptions are extended to include such moduli of uniform integrability for $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]$ and $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$, then we can effectively construct our modulus $\mu$ as follows:
Lemma 3.21.
@@
Suppose that $u$ and $x_{0}$ are chosen to be constant, that $K>0$ is such that $\left\lVert x_{0}-p\right\rVert,\left\lVert u-p\right\rVert<K$, and in addition both $\sum_{i=0}^{\infty}\left\lVert\xi_{i}\right\rVert$ and $\sum_{i=0}^{\infty}\left\lVert\delta_{i}\right\rVert$ are integrable with moduli of absolute continuity $\mu_{1}$ and $\mu_{2}$ respectively. Then a modulus of uniform integrability for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ is given by
$\mu(\varepsilon):=\min\left\{\frac{\varepsilon}{8K},\mu_{1}\left(\frac{\varepsilon}{8}\right),\mu_{2}\left(\frac{\varepsilon}{8}\right)\right\}.$
Proof.
By Lemma 3.9 we have $\left\lVert Uy_{n}-y_{n}\right\rVert\leq Y$ pointwise everywhere for all $n\in\mathbb{N}$, where $Y$ is defined as in Lemma 3.9, and so in particular we then have
$\left\lVert Uy_{n}-y_{n}\right\rVert\leq 2\left(2K+\sum_{i=0}^{\infty}\left(\left\lVert\xi_{n}\right\rVert+\left\lVert\delta_{n}\right\rVert\right)\right).$
for any $n\in\mathbb{N}$. Thus for any $A\in\mathcal{F}$ and $n\in\mathbb{N}$ it follow that
$\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert 1_{A}]\leq 4K\mathbb{P}(A)+2\left(\sum_{i=0}^{\infty}\left(\mathbb{E}[\left\lVert\xi_{i}\right\rVert 1_{A}]+\mathbb{E}[\left\lVert\delta_{i}\right\rVert 1_{A}]\right)\right)$
and so the result follows by definition of $\mu_{1}$ and $\mu_{2}$. ∎
While such moduli $\mu_{1},\mu_{2}$ always exist, the question of how these and related moduli of uniform integrability can actually be constructed in concrete situations still remains. First, we observe that in scenarios where $(\left\lVert Uy_{n}-y_{n}\right\rVert)$, or its $L_{1}$-bound $Y$ or the series over the errors enjoy stronger properties, such as higher moment conditions, corresponding moduli can be given in a straightforward manner and moreover assume a very simple form.
For that, we first consider the following general result:
Lemma 3.22.
Let $(X_{n})$ be a sequence of random variables such that $\sup_{n\in\mathbb{N}}\mathbb{E}[g(|X_{n}|)]<\infty$ for some measurable supercoercive $g:[0,\infty)\to[0,\infty)$, i.e. $g(x)/x\to\infty$ as $x\to\infty$. Then $(X_{n})$ is uniformly integrable and $\mu$ defined by
$\mu(\varepsilon):=\frac{\varepsilon}{2}\left(\kappa\left(\frac{2K}{\varepsilon}\right)\right)^{-1}$
is a corresponding modulus in the sense of Definition 3.16, where $\sup_{n\in\mathbb{N}}\mathbb{E}[g(|X_{n}|)]<K$ and $\kappa:(0,\infty)\to(0,\infty)$ is a rate of divergence for $g(x)/x\to\infty$, that is
$\forall a>0\,\forall x\geq\kappa(a)\left(\frac{g(x)}{x}\geq a\right).$
Proof.
Uniform integrability of $(X_{n})$ is a standard fact [41] (see also [17, Theorem 6.19]), and we simply need to verify the moduli. Fix $\varepsilon>0$, $n\in\mathbb{N}$ and $A\in\mathcal{F}$. Note that for any $a>0$, we have
$\mathbb{E}[|X_{n}|1_{A}]$ $\leq\mathbb{E}[|X_{n}|1_{A\cap(|X_{n}|\leq\kappa(a))}]+\mathbb{E}[|X_{n}|1_{A\cap(|X_{n}|>\kappa(a))}]$
$\leq\kappa(a)\mathbb{P}(A)+\mathbb{E}[|X_{n}|1_{A\cap(|X_{n}|>\kappa(a))}].$
Using now that $x>\kappa(a)$ implies $g(x)/a\geq x$, we have
$\mathbb{E}[|X_{n}|1_{A}]\leq\kappa(a)\mathbb{P}(A)+a^{-1}\mathbb{E}[g(|X_{n}|)]<\kappa(a)\mathbb{P}(A)+a^{-1}K.$
Setting $a:=2K/\varepsilon$ yields $\mathbb{E}[|X_{n}|1_{A}]\leq\kappa(2K/\varepsilon)\mathbb{P}(A)+\varepsilon/2$, so that if $\mathbb{P}(A)\leq\mu(\varepsilon)$ one has $\mathbb{E}[|X_{n}|1_{A}]\leq\varepsilon$. ∎
The above lemma is a quantitative variant of the fundamental de la Vallée-Poussin theorem [41] (see again also [17, Theorem 6.19]), and the existence of such a supercoercive function in fact characterizes uniformly integrable sequences of random variables. In particular, an immediate consequence is the following result for higher moment conditions:
Lemma 3.23.
Let $(X_{n})$ be a sequence of random variables such that $\sup_{n\in\mathbb{N}}\mathbb{E}[|X_{n}|^{p}]<K$ for some $K>0$ and $p>1$. Then $(X_{n})$ is uniformly integrable and $\mu$ defined by
$\mu(\varepsilon):=\frac{\varepsilon}{2}\left(\frac{\varepsilon}{2K}\right)^{1/(p-1)}$
is a corresponding modulus in the sense of Definition 3.16.
Lemmas 3.22 and 3.23 and can be applied directly to $(\left\lVert Uy_{n}-y_{n}\right\rVert)$, and of course in the special case that $X_{n}:=Y$ become simpler results on absolute continuity that can be instantiated to produce moduli of continuity on $Y$, and thus a modulus of integrability for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$. However, when reducing the problem to continuity properties of $\sum_{i=0}^{\infty}\left\lVert\xi_{i}\right\rVert$ and $\sum_{i=0}^{\infty}\left\lVert\delta_{i}\right\rVert$, as possible under the assumptions of this section as shown in Lemma 3.21, we conjecture that stronger assumptions such as those of Lemma 3.23 are not even required for the quantitative result, and that for concrete instantiations of the error terms (via e.g. minibatching as discussed in Section 5.1), moduli of integrability for the sums $\sum_{i=0}^{\infty}\left\lVert\xi_{i}\right\rVert$ and $\sum_{i=0}^{\infty}\left\lVert\delta_{i}\right\rVert$ can be calculated explicitly, exploiting the fact that we have concrete knowledge of the distribution of the errors. However, we do not give further details here.